# Are EU low-carbon structural funds efficient in reducing emissions?

### v1

@@ Abstract
This paper investigates the effectiveness of the “low-carbon economy” expenditures from European Structural and Investment Funds in fostering reductions in greenhouse gas emissions within European regions, focusing on the 2007-2013 and 2014-2020 programme periods. By decomposing emissions time series into trend and cycle components and considering them within a panel data framework, our research highlights that the impacts of low-carbon economy expenditures vary, qualitatively and quantitatively, with the targeted regions’ development level. We find significant emissions reductions in developed and transition regions yet less favourable outcomes in less developed areas. Further analysis into specific greenhouse gas emissions types (CO2, CH4, and N2O) reveals inconsistent impacts, underscoring the complexity of achieving emissions reductions. Our findings emphasise the need for tailored environmental strategies that accommodate the economic disparities of regions in the European Union.
Keywords: Greenhouse gas emissions; Emission decomposition; Low-carbon economy; European Structural and Investment Funds; Regional development
@@ 1 Introduction
The European Union (EU) has the objective of achieving climate-neutrality by 2050. This implies large-scale investments in the energy transition (McCollum et al., 2018; Klaassen and Steffen, 2023). According to the European Commission, energy investments in the EU will have to reach €396 billion per year from 2021 to 2030 and €520-575 billion per year in the subsequent decades until 2050 (European Parliament, 2023). A substantial share of these ought to be public investments that foster the development of infrastructure, the deployment of technological innovations, and the scale-up of private capital (Carraro et al., 2012; Owen et al., 2018; Meckling et al., 2022). Accordingly, the EU has enacted climate spending targets amounting to 20 of its 2014-2020 budget and 30 of its 2021-2027 budget that ought to be devoted to climate action111See European Court of Auditors (2022), which also emphasises that climate spending over the 2014-2020 period has been over-estimated.. In this context, one can ask a necessary question about the efficiency of European climate policy: have these expenses effectively led to the reduction of greenhouse gas (GHG) emissions?
@@
Consequently, we use the total amount spent on the low-carbon economy objective in the last two budgeting periods (2007-2013 and 2014-2020) as a measure of public climate investment. We then employ an econometric panel data model to investigate the influence of these investments on GHG emissions in European regions between 2007 and 2022 at the NUTS 2 level. Our methodology involves decomposing the emission time series of each region into trend and cycle components. This strategy gives us deeper insights into the nexus between economic growth and environmental sustainability. Due to the heterogeneity of European regions, it is clear that some have significantly advanced in terms of sustainability, achieving various degrees of decoupling, whereas many others have yet to do so. This disparity must be accounted for in the analysis to determine such expenditures’ short-term and long-term effects. Our analysis also accounts for the direct impacts of policy by considering the OECD Environmental Policy Stringency (EPS) index (OECD, 2016) and mean regional political opinions as covariates.
We find heterogeneous impacts of EU structural fund investments in low-carbon economies on GHG emissions reductions across European regions, with findings highlighting the variability of these effects by regional development levels, investment periods, and emission types. In developed and transition regions, low-carbon investments significantly lowered the long-term emission trends, in both per capita and relative to GDP, across the 2007-2013 and 2014-2020 programs. However, in the least developed regions, “low-carbon” investments had a positive (and significant) impact on emissions. Additionally, the effectiveness of reducing specific gas emissions, including CO2, CH4, and N2O, exhibits variability, indicating a complex interaction between investment efficiency and pollutant type. Our analysis also discerns significant but divergent impacts on the cyclical emissions components in transition economies between the two program periods. Hence, our analysis suggests a partial inefficiency of low-carbon public investments in Europe. It calls for a refined spatial and temporal allocation of these investments.
The heterogeneous impacts of the climate action efforts derived from the ESI funds can be attributed to several critical factors. These factors include the inherent complexities of reducing emissions, the intrinsic differences among European regions in terms of capabilities and resource endowments, and the relevance of development paths, as suggested by the Kuznets environmental curve (Bataille et al., 2021; Saraji and Streimikiene, 2023). For instance, it is possible that any additional funding may increase energy consumption and GHG emissions for investments below a certain level. This is rooted in how investments are targeted and their immediate impacts, which might not always be aligned with long-term sustainability goals. In many developing regions, the primary focus of initial investments may be on establishing basic infrastructure and boosting economic output quickly to meet immediate developmental needs. Additionally, low-carbon economy investments typically require substantial up-front costs and the deployment of heavy infrastructure, which can drive significant energy consumption and emissions. The transition to sustainable energy systems includes upgrading grids, developing renewable energy sources, and improving energy resilience. Though essential in the long term, these investments can increase short-term emissions due to the intensive energy and resources required during their implementation phase. On another front, there is a risk that investments have been used inefficiently or inappropriately, for example, as a countercyclical buffer to stabilise the economy during recessions or by investing in large infrastructure projects (aka “white elephants”) that do not generate the expected impacts. Arguments like these could align with the heterogeneous effects of public investment on climate outcomes.
@@
Therefore, the differential impacts on GHG emissions have their origin in the cyclical nature of economic activities. Recent literature offers valuable insights into the interplay between environmental impacts and the business cycle. Doda (2014) finds that CO2 emissions increase during economic upturns, exhibit greater volatility compared to GDP, and show a procyclicality that is more pronounced in higher GDP per capita economies, suggesting wealthier economies have more stable emission patterns through economic cycles. Sheldon (2017) finds an asymmetrical response of emissions to economic changes attributed to the timing of capital investments and technological advancements, indicating the significant role of corporate decisions in environmental impacts. Moreover, Klarl (2020) finds CO2 emissions’ elasticity to GDP changes to be significantly higher during recessions in the United States, showcasing the dynamic elasticity of emissions in response to economic conditions.
@@
In addition, climate and energy policy tools have proven some degree of effectiveness, but their success often hinges on countries’ stringency and the complementarity among strategies. Diakoulaki and Mandaraka (2007) find that despite considerable efforts, the expected acceleration in decoupling post-Kyoto Protocol was not uniformly achieved. For Europe, Naqvi and Zwickl (2017) and Ivanova et al. (2017) also observe significant disparities in decoupling across countries, sectors, pollutants, and regions. Papież et al. (2022, 2021) further report mixed results in the EU’s long-term decoupling efforts, pointing to disparities between older and newer member states and emphasising the role of energy intensity and structure in emissions reductions. For OECD countries, Chen et al. (2018), focusing on the key factors affecting CO2 emissions, conclude that while reductions in energy intensity are associated with decreased CO2 emissions, economic growth inevitably tends to increase CO2 emissions. This implies that emissions reductions are hard to accomplish unless economic growth is accompanied by significant improvements in energy efficiency and a shift towards less carbon-intensive energy sources.
Based on the literature discussed above, our paper contributes to the analysis of sustainable development and economic policy in European regions. By analysing how low-carbon investments from structural funds impact emissions reductions within business cycle fluctuations, we contribute to the understanding of decoupling economic growth from environmental degradation. Our results emphasise the importance of tailored investment strategies in achieving sustainable development goals, with a special focus on less developed regions to mitigate potential increases in emissions as they grow economically.
@@
The remainder of this study is structured as follows. Section 2 discusses the relevant strands of literature and the conceptual framework of our study and situates our contribution. Section 3 presents the data and methods. Section 4 presents the empirical results of the econometric estimations. Finally, Section 5 concludes by discussing the policy implications of our results.
@@ 2 Decoupling emissions: related literature
Our study builds on the extensive literature analysing the relationship between economic growth and GHG emissions. The Environmental Kuznets Hypothesis indicates that there is a connection between a country’s income level and the quality of its environment: it can be said that an increase in wealth generates greater demand from the population for a cleaner and healthier environment. This demand, in turn, drives governments to implement and enforce stricter environmental laws and regulations to protect the environment. While wealthier countries tend to have the resources and political will to adopt and enforce these effective environmental policies, regions with more constrained economic resources may find it more challenging to establish and maintain high environmental standards, resulting in more significant environmental challenges (O’Connor and Turnham, 1992; Doyle and Stiglitz, 2014).
@@
Certainly, economic development can lead to improved environmental outcomes by decoupling emissions from economic output. This is central to sustainable development and indicates that an economy can grow without a proportional increase in environmental degradation. However, the process is influenced by a multitude of factors, not only including income levels but also other aspects such as consumption patterns and regional characteristics. Shuai et al. (2017) analysed the country-by-country turning points of the carbon Kuznets curve, finding a positive correlation between a country’s income level and its likelihood to fit the Kuznets hypothesis, with higher income economies achieving more quick tipping points than their poorer counterparts. While economic development level is linked to faster environmental improvement transitions, other variables also play a relevant role. Ivanova et al. (2017) analyses household consumption across 177 regions in 27 EU countries, finding wide variations in per capita GHG emissions across regions (0.6 to 6.5 tCO2e/cap in 2007). Regarding the drivers behind carbon footprints, the authors identified the relevance of household size, whether an area is urban or rural, the population’s education level, expenditure patterns, local temperature, resource availability, and the carbon intensity of the local electricity mix.
@@
Yet a significant strand of literature doubts that the downward-sloping relationships between higher-income countries and pollution levels – which might imply a decoupling of the economic activity with emissions – is enough to face climate change efficiently. Certainly, there are concerns about the possible shift in the production of pollution-intensive goods from developed to developing countries with less stringent environmental regulations (Hertwich and Peters, 2009; Levinson, 2023). This would weaken the Kuznets hypothesis since developed regions may reduce their at-home pollution levels by offshoring the production of environmentally harmful goods to nations where environmental regulations are less strict (Fischer and Heutel, 2013). While this points out that high-income countries should also look into reducing aggregate demand as a strategy for significant emissions reductions (Knight and Schor, 2014), it also requires the implementation of agreements at different scales: regional and global (Oberthür and Ott, 1999).
@@
A growing body of evidence indicates that such policies and international agreements are increasingly impacting emission trends, as reflected in the economic practices within countries, contributing to our understanding of the decoupling mechanism (Fischer and Heutel, 2013; Hildén et al., 2014; Haberl et al., 2020). Decoupling analyses aim to categorise the patterns of how emissions respond to economic growth – whether they increase at a slower rate, decrease, or stabilise as the economy grows (relative and absolute decoupling) or increase faster than the economy (coupling).
@@
Cohen et al. (2018) provide valuable quantitative analysis that helps recognise how major emitting countries are progressing in decoupling emissions. Using information from OECD countries for 2001-2015, Chen et al. (2018) analyse how economic growth can be decoupled from CO2 emissions, identifying three key forms: recessive, where emissions drop due to efficiency gains or energy mix shifts, not economic growth; weak negative, where efficiency improvements do not fully offset emissions from economic expansion; and strong, where economic and population growth do not lead to higher emissions, thanks to effective policies, technological progress, and structural economic changes.
@@
Building on this understanding of the decoupling patterns, it is imperative to analyse emissions in the context of business cycles due to their substantial implications for environmental policy and economic management, particularly regarding the relationship between economic activity, regulation, emissions and energy consumption (Fischer and Heutel, 2013). In fact, emissions tend to be procyclical, increasing during economic booms and decreasing during recessions (Heutel, 2012). A fundamental question relates to the causal relationship between economic output and energy consumption, especially targeting business cycle fluctuations (Thoma, 2004; Narayan et al., 2011). Energy consumption notably responds to changes in industrial production and macroeconomic conditions, reinforcing the concept of the procyclicality of emissions relative to GDP observed in broader research. For instance, Thoma (2004) demonstrates that changes in macroeconomic conditions cause significant changes in electrical usage, particularly in the commercial and industrial sectors.
@@
Therefore, the link between economic growth, energy consumption, and GHG emissions is clearly established in the short term. For example, Peters et al. (2012) claim that economic downturns typically lead to a reduction in energy-intensive activities and lower CO2 emissions. Notwithstanding, the authors discuss that the 2008-2009 financial crisis saw a quick dip in GDP followed by a swift rebound in emissions by 2010, driven by falling energy prices, government economic incentives, and prior economic growth in developing countries, quickly returning to high emission levels post-crisis. Similarly, Doda (2014) finds that CO2 emissions are procyclical and more pronounced in countries with higher GDP per capita. Hence, the procyclicality also implies that economic growth without specific interventions often leads to increased emissions. Another fact is that emissions exhibit greater volatility compared to GDP, responding more dramatically to economic changes, and the extent to which emissions volatility exceeds GDP volatility decreases with the income level of the economies (Doda, 2014). This implies that more stable emission patterns through economic cycles in developed regions could be linked to more efficient energy use, a greater share of services in the economy, and potentially more effective environmental regulations and policies. In addition, emissions decrease more significantly during economic contractions than during expansions, suggesting an elastic response to economic downturns and inelasticity to economic growth (Sheldon, 2017).
@@
Likewise, Klarl (2020) examines the fluctuation of CO2 emissions’ elasticity to GDP in the United States from 1973 to 2015, revealing that this elasticity is not constant and varies with economic conditions. He showed that emissions respond more sensitively to GDP changes during recessions than periods of economic normality, with elasticity often exceeding one in downturns. This indicates that CO2 emissions decrease more rapidly than economic output in recessions, highlighting the impact of economic cycles on environmental outcomes. The finding that emissions exhibit greater sensitivity during economic downturns suggests that the capacity of an economy to influence environmental outcomes can be significantly conditioned by its current economic state.
@@
In a similar vein, Naqvi (2021) finds that, on average, EU regions have managed to decouple emissions from economic output, with the most significant emission reductions occurring before the 2008 financial crisis. After 2008, the trend towards decoupling weakened, with instances of re-coupling between emissions and economic output observed. The author uses the EPS index, finding that stronger environmental policies effectively reduce emissions. However, the effectiveness varies significantly across different types of emissions and is also influenced by the income levels of the regions. Similarly, with the aim of monitoring the effectiveness of the EU’s environmental policies, Delgado Rodríguez et al. (2018) developed the EU-cyclical environmental performance index by applying dynamic factor analysis over the period 1950 to 2012. This index demonstrates that the implementation of stricter emission targets, initiated particularly since the adoption of the 5th Environmental Action Programme (EAP), has effectively contributed to restraining the growth of CO2 emissions across the EU. Furthermore, the index reveals heterogeneity among the member states, distinguishing between countries that are leading in emission linkage – thus enhancing their alignment with EU environmental standards – and those lagging behind, which may require additional incentives or policy measures to improve compliance with EU regulations. In addition, Papież et al. (2022) evaluate the EU’s energy policy effectiveness in achieving long-term decoupling and reports mixed findings. The study notes progress in detaching economic growth from emissions under the EU ETS and production-based accounting in both EU-15 and new member countries. However, they find no advancement in separating economic growth from consumption-based accounting emissions in either group.
@@
In summary, environmental policies must confront the broader economy, influenced by external macroeconomic shocks and internal technological changes. Under Kuznets’ hypothesis, decoupling might appear problematic since countries would only begin to improve their environmental performance after reaching a certain level of wealth, implying a dependency between economic growth and environmental improvement. Therefore, seeking economic development without the corresponding environmental degradation is a great challenge for any policy framework.

### v2

@@ Abstract
We investigate the effectiveness of low-carbon expenditures from the European Structural and Investment Funds in reducing greenhouse gas emissions across EU regions. Using trend and cycle decomposition of two indicators – per capita emissions and emissions intensity – we find highly heterogeneous effects. In less developed regions, investments are linked to long-term increases in per capita emissions, while in transition and developed regions, effects are weak or not significant. Instrumental variable estimates confirm this pattern. When disaggregated by gas type, results remain inconsistent. Overall, our findings highlight the limited climate effectiveness of EU funding and the need for region-specific strategies.
Keywords: Greenhouse gas emissions; Green transition; Climate action; Low-carbon economy; European Structural and Investment Funds; Regional development
@@ 1 Introduction
The European Union (EU) has the objective of achieving climate neutrality by 2050. This implies large-scale investments in the energy transition (McCollum et al., 2018; Klaassen and Steffen, 2023). According to the European Commission, energy investments in the EU will have to reach €396 billion per year from 2021 to 2030 and €520-575 billion per year in the subsequent decades until 2050 (European Parliament, 2023). A substantial share of these ought to be public investments that foster the development of infrastructure, the deployment of technological innovations, and the scale-up of private capital (Carraro et al., 2012; Owen et al., 2018; Meckling et al., 2022). Accordingly, the EU has enacted climate spending targets amounting to 20 of its 2014-2020 budget and 30 of its 2021-2027 budget that ought to be devoted to climate action111See European Court of Auditors (2022), which also emphasises that climate spending over the 2014-2020 period has been over-estimated.. Therefore, it is reasonable to question the efficiency of European climate policy to determine if these expenses effectively lead to reducing greenhouse gas (GHG) emissions.
@@
Consequently, we use the total amount spent on the low-carbon economy objective for the 2007-2013 and 2014-2020 budgeting periods as a measure of public climate investment. We then employ an econometric panel data model to investigate the influence of these investments on GHG emissions in European regions between 2007 and 2022 at the NUTS 2 level. Our analysis centres on two emission indicators: per capita emissions and emissions intensity (relative to regional gross domestic product (GDP)). We focus on the relationship between economic growth and environmental sustainability, and, to that end, we decompose the emission time series of each region into trend and cycle components. This allows us to determine expenditures’ short-term and long-term effects, considering that some regions have significantly advanced in terms of sustainability – achieving a given degree of decoupling – whereas many others have yet to do so. Our analysis also accounts for the direct impacts of policy by considering the OECD Environmental Policy Stringency (EPS) index (OECD, 2016) and average regional political opinions as covariates.
We find heterogeneous effects of low-carbon expenditures on GHG emissions across European regions. These impacts vary depending on the regional development level, investment period, and emission indicator. In the least developed regions, low-carbon expenditures are consistently associated with increases in long-term emissions per capita and, to a lesser extent, in emissions intensity. In contrast, results for developed and transition regions do not reveal significant reductions in long-term emissions trends, although there is some evidence of short-term mitigating effects, particularly for the 2007–2013 programme.
To address potential endogeneity, we implement a two-stage least squares (2SLS) model using lagged structural indicators and regional development categories as instruments. The instrumental variables estimates reinforce the main findings: low-carbon investments have heterogeneous effects across regions, with increases in per capita emissions in less developed regions and short-term reductions in developed regions. In transition regions, distinct patterns emerge that ultimately underscore the importance of the regional context.
When disaggregating emissions by gas type, we consistently find marked regional heterogeneity. The results reveal no significant reductions in emissions for developed regions, while in less developed regions, emissions – especially CO2 and N2O per capita – increase significantly. Effects in transition regions are mixed. Short-term dynamics also show variation by gas and region: investments under the 2007–2013 programme help reduce emissions volatility, particularly for CO2 in less developed and developed regions, and for N2O in transition regions.
Overall, our findings point to important challenges in the effectiveness and coherence of current EU climate investment strategies, particularly those guided by the ESI funds, in supporting long-term decarbonisation in less developed regions. However, rather than signalling a failure of climate investment itself, the uneven performance may reflect the persistent structural regional disparities, including differences in institutional capacity, infrastructure, and development trajectories, as suggested by the Kuznets environmental curve (Bataille et al., 2021; Saraji and Streimikiene, 2023). Investments’ immediate impacts are not necessarily always aligned with long-term sustainability goals. For instance, it could be that investments below a certain level may increase energy consumption and raise GHG emissions. In many developing regions, investments are focused on establishing basic infrastructure to boost economic output and meet developmental needs.
In addition, low-carbon economy investments might require substantial up-front costs (deploying heavy infrastructure), and sustainable energy systems demand upgrading grids, developing renewable energy sources, and improving energy resilience. Though essential in the long term, these investments can increase short-term emissions due to the intensive energy and resources required during their implementation phase. On another front, investments can be misused, for example, as a countercyclical buffer to stabilise the economy during recessions or in implementing large and inefficient infrastructures (aka “white elephants”).
@@
While arguments behind the heterogeneous impacts vary across regions, they are certainly shaped by economic conditions, particularly the economic cycle. A recent strand of literature offers insights into the interplay between environmental impacts and the business cycle. Doda (2014) finds that CO2 emissions increase during economic upturns, exhibit greater volatility compared to GDP, and show a procyclicality that is more pronounced in higher GDP per capita economies, suggesting that wealthier economies have more consistent emission patterns through economic cycles. Sheldon (2017) finds an asymmetrical response of emissions to the business cycle, which is attributed to the timing of capital investments and technological advancements. Moreover, Klarl (2020) finds CO2 emissions’ elasticity to GDP changes to be significantly higher during recessions in the United States. These results demonstrate the critical role of investment decisions in shaping environmental impacts.
@@
In addition, climate and energy policy tools have shown some degree of effectiveness, however, their success often relies on countries’ stringency and the complementarity among strategies. Diakoulaki and Mandaraka (2007) find that despite considerable efforts, the expected acceleration in decoupling post-Kyoto Protocol was not uniformly achieved. For Europe, Naqvi and Zwickl (2017) and Ivanova et al. (2017) find significant disparities in decoupling across countries, sectors, pollutants, and regions. Papież et al. (2022, 2021) further report mixed results in the EU’s long-term decoupling efforts, pointing to disparities between older and newer member states and emphasising the role of energy intensity and structure in emissions reductions.
In light of these findings, our paper contributes to the understanding of decoupling economic growth from emissions by analysing how low-carbon investments from the ESI funds impact emissions reductions within the business cycle.
@@
The remainder of this study is structured as follows. Section 2 discusses the relevant literature and the conceptual framework of our study and situates our contribution. Section 3 presents the data and methods. Section 4 presents the empirical results of the econometric estimations. Section 5 discusses the results and provides policy implications. Finally, Section 6 concludes.
@@ 2 Decoupling emissions: related literature
Our study builds on the extensive literature on the connection between economic growth and GHG emissions. Drawing from the Environmental Kuznets Hypothesis, it is assumed that an increase in wealth generates greater demand from the population for a cleaner and healthier environment. This demand drives governments to implement stricter environmental regulations to protect the environment. In this way, while wealthier countries tend to have the resources and political will to adopt these policies, regions with more constrained economic resources may find it more challenging to establish high environmental standards (O’Connor and Turnham, 1992; Doyle and Stiglitz, 2014).
@@
Certainly, economic development can lead to improved environmental outcomes by decoupling emissions from economic output. This is central to sustainable development and indicates that an economy can grow without a proportional increase in environmental degradation. However, the process is influenced by a multitude of factors, not only including income levels but also other aspects such as consumption patterns and regional characteristics. Shuai et al. (2017) analyse the country-by-country turning points of the carbon Kuznets curve, finding a positive correlation between a country’s income level and its likelihood to fit the Kuznets hypothesis, with higher income economies achieving quicker tipping points than their poorer counterparts. While economic development level is linked to faster environmental improvement transitions, other variables such as consumption patterns and regional characteristics also play a relevant role. Ivanova et al. (2017) analyse household consumption across 177 regions in 27 EU countries, finding wide variations in per capita GHG emissions across regions (0.6 to 6.5 tCO2e/cap in 2007). Regarding the drivers behind carbon footprints, the authors identified the relevance of household size, whether an area is urban or rural, the population’s education level, expenditure patterns, local temperature, resource availability, and the carbon intensity of the local electricity mix.
@@
Yet a significant strand of literature doubts that the downward-sloping relationships between higher-income countries and pollution levels – which might imply a decoupling of economic activity with emissions – are enough to face climate change efficiently. Certainly, there are concerns about the possible shift in the production of pollution-intensive goods from developed to developing countries with less stringent environmental regulations (Hertwich and Peters, 2009; Levinson, 2023). This would weaken the Kuznets hypothesis since developed regions may reduce their at-home pollution levels by offshoring the production of environmentally harmful goods to nations where environmental regulations are less strict (Fischer and Heutel, 2013). While this points out that high-income countries should also look into reducing aggregate demand as a strategy for significant emissions reductions (Knight and Schor, 2014), it also requires the implementation of agreements at different scales: regional and global (Oberthür and Ott, 1999).
@@
In this context, a growing body of literature indicates that international environmental agreements and climate policies are increasingly impacting emission trends, as reflected in the economic practices within countries, contributing to our understanding of the decoupling mechanism (Fischer and Heutel, 2013; Hildén et al., 2014; Haberl et al., 2020). Decoupling analyses aim to categorise the patterns of how emissions respond to economic growth – whether they increase at a slower rate, decrease, or stabilise as the economy grows (relative and absolute decoupling) or increase faster than the economy (coupling).
@@
Cohen et al. (2018) provide valuable quantitative analysis that helps recognise how major emitting countries are progressing in decoupling emissions. Using information from OECD countries for 2001-2015, Chen et al. (2018) analyse how economic growth can be decoupled from CO2 emissions, identifying three key forms: recessive, where emissions drop due to efficiency gains or energy mix shifts, not from economic growth; weak negative, where efficiency improvements do not fully offset emissions from economic expansion; and strong, where economic and population growth do not lead to higher emissions, thanks to effective policies, technological progress, and structural economic changes.
@@
Building on this understanding of the decoupling patterns, it is imperative to analyse emissions in the context of business cycles due to their substantial implications for environmental policy and economic management, particularly regarding the relationship between economic activity, regulation, emissions and energy consumption (Fischer and Heutel, 2013). In fact, emissions tend to be procyclical, increasing during economic booms and decreasing during recessions (Heutel, 2012). A fundamental question relates to the causal relationship between economic output and energy consumption, especially targeting business cycle fluctuations (Narayan et al., 2011). Thoma (2004) shows that changes in macroeconomic conditions cause significant changes in electrical usage, particularly in the commercial and industrial sectors. These energy consumption responses reinforce the procyclicality of emissions relative to GDP.
Therefore, the connection among economic growth, energy consumption, and GHG emissions is well established in the short term. For example, Peters et al. (2012) claim that economic downturns typically lead to a reduction in energy-intensive activities and lower CO2 emissions. Notwithstanding, the authors discuss that the 2008-2009 financial crisis saw a quick dip in GDP followed by a swift rebound in emissions by 2010, driven by falling energy prices, government economic incentives, and prior economic growth in developing countries, quickly returning to high emission levels post-crisis.
@@
Similarly, Doda (2014) finds that CO2 emissions are procyclical and more pronounced in countries with higher GDP per capita. Hence, the procyclicality also implies that economic growth without specific interventions often leads to increased emissions. Another fact is that emissions exhibit greater volatility compared to GDP, responding more dramatically to economic changes, and the extent to which emissions volatility exceeds GDP volatility decreases with the income level of the economies (Doda, 2014). These patterns are also observed at the subnational level: Dueñas and Mandel (2025), analysing NUTS 2 regions in Europe, further show that CO2 emissions growth rates follow an asymmetric and fat-tailed distribution and, similarly, that emissions volatility scales inversely with economic size of regions. This implies that more stable emission patterns through economic cycles in developed regions could be linked to more efficient energy use, a greater share of services in the economy, and potentially more effective environmental regulations and policies. In addition, emissions decrease more significantly during economic contractions than during expansions, suggesting an elastic response to economic downturns and inelasticity to economic growth (Sheldon, 2017).
@@
Likewise, Klarl (2020) examines the fluctuation of CO2 emissions’ elasticity to GDP in the United States from 1973 to 2015, revealing that this elasticity is not constant and varies with economic conditions. He showed that emissions respond more sensitively to GDP changes during recessions than periods of economic normality, with elasticity often exceeding one in downturns. This indicates that CO2 emissions decrease more rapidly than economic output in recessions and also suggests that the capacity of an economy to influence environmental outcomes can be significantly conditioned by its current economic state.
Thus, the literature has established the relevance of economic policy in addressing both the cyclical sensitivity of emissions to economic activity and the long-term decoupling of emissions from economic growth. This is of particular interest in the context of the European Union, where studies have started to examine the role of policy in shaping both short-term dynamics and long-term environmental trends.
@@
For example, Naqvi (2021) finds that, on average, EU regions have managed to decouple emissions from economic output, with the most significant emission reductions occurring before the 2008 financial crisis. After 2008, the trend towards decoupling weakened, with instances of re-coupling between emissions and economic output observed. In addition, the author finds that environmental policy effectiveness varies significantly across different types of emissions and is influenced by the income levels of the regions. Similarly, with the aim of monitoring the effectiveness of the EU’s environmental policies, Delgado Rodríguez et al. (2018) developed the EU-cyclical environmental performance index by applying dynamic factor analysis over the period 1950 to 2012. This index shows that the implementation of stricter emission targets, initiated particularly since the adoption of the 5th Environmental Action Programme (EAP), has effectively contributed to restraining the growth of CO2 emissions across the EU. Furthermore, the index is substantially heterogeneous among the member states, distinguishing between countries that lead to emission linkage and those lagging behind. In addition, Papież et al. (2022) evaluate the EU’s energy policy effectiveness in achieving long-term decoupling and reports mixed findings. The study notes progress in detaching economic growth from emissions under the EU ETS and production-based accounting, but they find no advancement in separating economic growth from consumption-based accounting emissions in either group.
More directly related to the impacts of the EU Cohesion Policy, much of the literature has focused primarily on understanding its impact on regional development, specifically whether the funds have been effective in reducing regional disparities. For instance, very early on, Boldrin and Canova (2001) pointed out that these policies tend to be more redistributive than growth-oriented, reflecting political compromises over economic efficiency. However, investments in education and human capital, which have gained a significant portion of these funds, have shown medium-term positive impacts, emphasising the critical role of human capital development in fostering long-term regional growth and convergence (Rodríguez-Pose and Fratesi, 2004). Institutional quality plays a relevant role in the effectiveness of the funds. Evidence shown that local and regional governance directly affects economic growth and the effectiveness of structural funds (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014).
Although relatively few studies have explicitly focused on the effect of EU structural funds on reducing emissions, some anticipate that many of the same structural patterns that determine growth performance may also shape the trajectory of sustainable growth. Rodríguez-Pose and Bartalucci (2024) warn that while the green transition may benefit certain regions, it can also exacerbate existing territorial inequalities, particularly affecting less-developed regions. Nichifor et al. (2025) find that investments in renewables and environmental expenditures have positive effects on emissions reduction, but caution that these effects may exhibit diminishing returns beyond certain thresholds.
@@
In summary, in light of these findings, the effectiveness of the ESI funds in driving a low-carbon transition is conditional not only on structural factors such as institutional quality and regional development levels, but also on the phase of the economic cycle. Under Kuznets’ hypothesis, decoupling might appear problematic since regions would only begin to improve their environmental performance after reaching a certain level of wealth, implying a dependency between economic growth and environmental improvement. Therefore, seeking economic development without the corresponding environmental degradation is a great challenge for any policy framework.
Therefore, the central hypothesis of this paper is as follows:
Hypothesis: Low‑carbon economy expenditures from the European Structural and Investment Funds have heterogeneous impacts on GHG emissions across regions at different stages of development, influencing both the long‑run trend and the magnitude of the business‑cycle component of emissions.
To test this hypothesis, rather than focusing on total emissions, we concentrate on per capita emissions and emissions intensity (relative to regional GDP) . These indicators enhance comparability across regions by accounting for differences in population and economic scale, while also isolating environmental efficiency more effectively.